OpenStudy (anonymous):

Use the function f(x) = x^4 - 36 x^3 At what value(s) of x does f have a local maximum or a local minimum?

OpenStudy (anonymous):

x=0 at the slope of the curve but i couldnt find the other point

sam (.sam.):

dy/dx = 4x^3 - 108x^2 when dy/dx = 0, (That means maximum and minimum) 4x^3-108x^2 =0 x^2 (4x-108)=0 x=0 x=27

sam (.sam.):

dy/dx = 0 because when maximum or minimum point, the gradient is 0. and dy/dx is gradient.

OpenStudy (anonymous):

f'(x) = 4x^3 - 108x^2 = 0 -> 4x^2 ( x - 27) = 0 => x = 0, x = 27

OpenStudy (anonymous):

x = 0, y = 0 =>(0,0) Max x = 27, y = 27^3 ( 27 - 36) = -177147 ->( 27, -177147) min

Directrix (directrix):

x^3 (x-36) = 0 x^3 = 0 or x = 36 Triple root at 0 indicates that graph is tangent to x-axis --> neither max nor min Graph slices through x-axis at 36 Dominant term of x^4 indicates entry into Quad 2 and exit in Quad 1 Absolute minimum at x = 27 --> (27, - 177,147) (needs checking)